According to the Center for Disease Control, 41.5% of babies born in the U.S. in 2004 were still being breastfed at 6 months of age. If 4 children who were born
in the U.S. in 2004 are randomly selected, what is the probability that none of them were breastfed for at least 6 months.
I tried everything, I think the none of them is throwing me off.
I thought using P(at least one will not)-P(all will have been breastfed; but then that did not make sense.
Can you help?
Thank you,
Dawn

Answer by stanbon(57308)

According to the Center for Disease Control, 41.5% of babies born in the U.S. in 2004 were still being breastfed at 6 months of age.
------------------------------------------
If 4 children who were born in the U.S. in 2004 are randomly selected, what is the probability that none of them were breastfed for at least 6 months.
Ans: (1-0.415)^4 = 0.585^4 = 0.1171
------------------------------------------
Cheers,
Stan H.